Consider a standard deck of 52 cards. How many 5 card hands can be formed that include 2 hearts and 3 clubs
Answer by Edwin McCravy(20056)

13 hearts choose 2 = 13C2 = (13*12)/(2*1) = 78
For each of those 78 ways to choose 2 hearts, there are
13 clubs choose 3 = 13C3 = (13*12*11)/(3*2*1) = 286
ways to choose the 3 clubs.

Answer (13C2)(13C3) = (78)(286) = 22308 ways.

Edwin